HW4

Given the following information of the mortgage pool that backs a MPT, what is the regular scheduled payment in month 1 of the security? Use WAC as the mortgage rate and WAM as the number of periods for your calculations. Round your final answer to two decimals.

• 30 year FRM, fully amortizing, monthly payments

• Loans seasoned for 3 months before entering pool

• WAM: 357 

• WAC: 4%

• Servicer/Guarantee fee: 0.55%

• Starting pool balance: 250,342,967

• Prepayment assumption: 75% PSA


Given the following information of the mortgage pool that backs a MPT (same as Question 3,4), what is the total cash flow of the pool (i.e. to the issuer/SPV) in month 1 of this security? Round your final answer to two decimals.

• 30 year FRM, fully amortizing, monthly payments

• Loans seasoned for 3 months before entering pool

• WAM: 357 

• WAC: 4%

• Servicer/Guarantee fee: 0.55%

• Starting pool balance: 250,342,967

• Prepayment assumption: 75% PSA
Answer #1

a. To calculate the regular scheduled payment in month 1 of the security, we can follow the steps from the previous answer:

Step 1: Calculate the Monthly Mortgage Rate (MMR) based on the Weighted Average Coupon (WAC). MMR = WAC / 12 = 4% / 12 = 0.3333% (0.003333 as a decimal)

Step 2: Calculate the Monthly Payment (PMT): PMT = Starting pool balance * (MMR / (1 - (1 + MMR)^(-WAM))) PMT = 250,342,967 * (0.003333 / (1 - (1 + 0.003333)^(-357))) PMT ≈ 250,342,967 * 0.004450 PMT ≈ 1,113,712.59

So, the regular scheduled payment in month 1 of the security is approximately $1,113,712.59.

b. To calculate the total cash flow of the pool to the issuer/SPV in month 1, we need to consider the principal repayment and the Servicer/Guarantee fee:

Principal Repayment in month 1 = Starting pool balance * PSA Principal Repayment in month 1 = 250,342,967 * 0.75 Principal Repayment in month 1 = 187,757,225.25

Total Cash Flow = Monthly Payment - Principal Repayment + Servicer/Guarantee fee Total Cash Flow = 1,113,712.59 - 187,757,225.25 + (0.55% * 250,342,967) Total Cash Flow ≈ 1,113,712.59 - 187,757,225.25 + 1,377,386.82 Total Cash Flow ≈ -$185,266,126.84

The total cash flow to the issuer/SPV in month 1 of the security is approximately -$185,266,126.84 (a negative value indicates cash outflow).


answered by: Aratrika